Suppose a manufacture invented a computer chip in 1978 that had a computational speed of s. The company improves its chips so that every 3 years, the chip doubles in speed. What would the chip's speed have been for the year 2002? Wrute your solution in terms of s.
Answer by ankor@dixie-net.com(22740)

Suppose a manufacture invented a computer chip in 1978 that had a computational speed of s. The company improves its chips so that every 3 years, the chip doubles in speed. What would the chip's speed have been for the year 2002? Wrute your solution in terms of s.
:
How about this:
Years for this to happen: 2002 - 1978 = 24 years
Then the speed doubles 24/3 = 8 times in this period of time (t=8)
:
f(s) = s*(2^t)
f(s) = s*(2^8)
f(s) = 256s